# biblio.bib

### v1

@@ 2.1 Eigenspaces of the exchange (XX) Hamiltonian
We start by analyzing only the exchange part of the Hamiltonian (1), and considering it as our unperturbed mean-field component: a similar treatment can be carried out for the Z part alone, still in the case of constant coupling between every couple of sites.
^$\hat{\mathcal{H}}_{0}=J\sum_{i<j}(\hat{\mathcal{S}}_{i}^{x}\hat{\mathcal{S}}_{j}^{x}+\hat{\mathcal{S}}_{i}^{y}\hat{\mathcal{S}}_{j}^{y})=\frac{J}{2}\sum_{i<j}(\hat{\mathcal{S}}_{i}^{+}\hat{\mathcal{S}}_{j}^{-}+\hat{\mathcal{S}}_{i}^{-}\hat{\mathcal{S}}_{j}^{+})\,.$
We show below that ^$\hat{\mathcal{H}}_{0}$ can be rewritten as function of the total $z$-magnetization ^$\hat{\mathcal{M}}_{z}=\sum_{i}\hat{\mathcal{S}}_{i}^{z}$ and of the (Casimir invariant) total squared magnetization ^$\hat{\mathcal{M}}^{2}=\hat{\mathcal{M}}_{x}^{2}+\hat{\mathcal{M}}_{y}^{2}+\hat{\mathcal{M}}_{z}^{2}$, see [santos2016cooperative]. We then re-cast the operators $\hat{\mathcal{S}}_{i}^{x}\hat{\mathcal{S}}_{j}^{x}$ (equivalent for the $y$-component) in terms of ^$\hat{\mathcal{M}}_{z}$ and ^$\hat{\mathcal{M}}^{2}$ as:
^$\hat{\mathcal{M}}_{x}^{2}=\sum_{k}(\hat{\mathcal{S}}_{k}^{x})^{2}+2\sum_{i<j}\hat{\mathcal{S}}_{i}^{x}\hat{\mathcal{S}}_{j}^{x}\,\Leftrightarrow\,\sum_{i<j}\hat{\mathcal{S}}_{i}^{x}\hat{\mathcal{S}}_{j}^{x}=\frac{\hat{\mathcal{M}}_{x}^{2}-\sum_{k}(\hat{\mathcal{S}}_{k}^{x})^{2}}{2}\,.$
@@
^$\hat{\mathcal{H}}_{0}=\frac{J}{4}\Bigl{[}2(\hat{\mathcal{M}}^{2}-\hat{\mathcal{M}}_{z}^{2})-L\Bigr{]},$ (2)
with $L$ the system size. For reference, the respective computation made for the Z part only amounts to ^$\hat{\mathcal{H}}^{\prime}_{0}=J/2(\hat{\mathcal{M}}_{z}^{2}-L/4)$. The result in Eq. (2) implies that although the starting Hamiltonian is the sum of two-body interacting terms, the all-to-all connectivity allows for the appearance of collective behavior, where the system behaves as if it had a single, large spin [omanakuttan2023scrambling]. Labeling the quantum numbers associated to operators ^$\hat{\mathcal{M}}^{2}$ and ^$\hat{\mathcal{M}}_{z}$ with $l$ and $m$, respectively, we use the angular momenta composition rules for spin-$1/2$ to calculate the corresponding energy bands $E_{lm}$:
$l$ =$=\begin{cases}0,1,\dots,\frac{L}{2}\quad\text{if $L$ is even}\\ \frac{1}{2},\frac{3}{2},\dots,\frac{L}{2}\quad\text{if $L$ is odd}\\ \end{cases}$
$m$ $=-l,(-l+1),\dots,(l-1),l\quad\text{($2l+1$ values)}$)
@@
The degeneracy $f(l,L)$ of a certain value of $l$ for a given $L$, can be computed recursively with:
@@
$\text{deg}(E_{l|m|},L)=\begin{cases}2f(l,L)\quad\text{if $L$ is odd or $L$ is even and $m\neq 0$}\\ f(l,L)\quad\text{if $L$ is even and $m=0$}\\ \end{cases}\,.$
Figure 2: Symmetries and structure of the energy levels for the all-to-all connected XX Hamiltonian. $(a)$ Diagram for a system with $L=7$ sites and constant exchange coupling $J/2=1$ (unperturbed study). The vertical axis represents the exact values of the eigenvalues of ^$\hat{\mathcal{H}}_{0}$, while the numbers on the left of the horizontal lines account for the degeneracy of the corresponding level; different colors and symbols represent different subspaces corresponding to different number of excitations along the $z$-axis. Red triangles and lines: levels corresponding to 1 excitation. Blue lozenges and lines: levels referring to 2 excitations. Double-headed arrows : the three possible energy gaps described by Eq. (5): the dotted arrow represents the frequency that is not manifested in the z-magnetization spectrum. Green bars: remaining energy levels. $(b)$ The same diagram for the case $L=9$, with the red bars contoured by stars representing the levels with 4 excitations denoted by the quantum number $l$ on the right.
@@
We are now in a position to explain the local $z$-magnetization spectra in Fig. 1 in terms of the energy levels. Due to the $U(1)$ symmetry, the quantum number $m$ will be fixed during the time evolution. If we consider the case of one single excitation to start with, we have $m_{1}=(-L/2)+1$, hence only the values $l_{1}=(L/2)-1$ and $l_{2}=L/2$ are allowed, the corresponding bands having energies $E_{l_{1}m_{1}}=-{J}/2\doteq E_{1}^{(1)}$ with degeneracy $L-1$ and $E_{l_{2}m_{1}}={J(L-1)}/2\doteq E_{2}^{(1)}$ with degeneracy 1. Thus, the oscillation of the local $z$-magnetization signal occurs at angular frequency equal to the energy split between the two levels:
@@
We notice that the local $z$-magnetization spectra depend explicitly on the separation between the energy levels of the system.
We can use the same line of reasoning for the second case highlighted in Fig. 2$(a)$, with two excitations. Here, the quantum number $m$ is fixed to be $m_{2}=(-L/2)+2$, implying that $l$ can take values $l_{1}=(L/2)-2$, $l_{2}=(L/2)-1$ and $l_{3}=L/2$, reflecting in the energy levels as $E_{l_{1}m_{2}}=-J\doteq E_{1}^{(2)}$ with degeneracy $L(L-3)/2$; $E_{l_{2}m_{2}}={J(L-4)}/2\doteq E_{2}^{(2)}\,\text{with degeneracy $L-1$;}$ $E_{l_{3}m_{2}}=J(L-2)\doteq E_{3}^{(3)}\,\text{ with degeneracy 1.}$. The resulting energy gaps are:
@@
Focusing on the case of two excitations, we analyze in Fig. 3 the power spectrum of the local magnetization $\langle\hat{\mathcal{S}}_{v}^{z}\rangle(t)$. In Fig. 3(a), we observe the presence of two frequencies, corresponding to the gaps $\Delta E_{21}$ and $\Delta E_{32}$, while the third possible transition is absent. Even in the case depicted in Fig.1$(d)-(e)$, corresponding to four initial excitations or $m=-L/2+4$, we ascertain the presence of four out of the $4\dot{(}4-1)/2=6$ possible frequencies arising from the five levels having quantum number $l=L/2,L/2-1,L/2-2,L/2-3,L/2-4$; again, these four frequencies are exactly those coming from the difference between contiguous levels. The schematics for this situation is shown in Fig. 2$(b)$.
@@
where we have decomposed the initial state $|\psi_{0}\rangle$ temporal evolution in the Hamiltonian eigenbasis $\{|\phi_{n}\rangle\}$, with corresponding eigenvalues $\{E_{n}\}$ (the pedices here represent a generic quantum number). In our specific case, the eigenstates of the system Hamiltonian ^$\hat{\mathcal{H}}_{0}$ are those of the angular momentum $|l,m\rangle$, with $l$ and $m$ defined in Eq. 3, the initial state is a superposition of Pauli strings, and the operator $\hat{\mathcal{O}}$ is a local spin operator $\hat{\mathcal{S}}^{z}_{i}\doteq\hat{\mathcal{O}}_{0}^{(1)}$ at site $i$: this operator has the property of being the $q=0$ component of a spherical tensor of rank $r=1$, the other components being $\hat{\mathcal{O}}_{+1}^{(1)}=-c\hat{\mathcal{S}}^{+}_{i}$ and $\hat{\mathcal{O}}_{-1}^{(1)}=c\hat{\mathcal{S}}^{-}_{i}$ with $c$ a positive constant. The coefficients $\langle\psi_{0}|\phi_{m}\rangle=\langle\psi_{0}|\alpha;l,m\rangle$ and $\langle\phi_{n}|\psi_{0}\rangle=\langle\beta;l^{\prime},m^{\prime}|\psi_{0}\rangle$ are the Clebsch-Gordan coefficients between the eigenbasis and the Pauli basis, where $\alpha$ and $\beta$ denote the system-specific quantum numbers not related to angular momentum. The matrix elements of $\hat{\mathcal{O}}_{0}^{(1)}$ can be expressed thanks to the Wigner-Eckart theorem [sakurai1994] as:
$\langle\phi_{m}|\hat{\mathcal{O}}|\phi_{n}\rangle=\langle\alpha;l,m|\hat{\mathcal{O}}_{0}^{(1)}|\beta;l^{\prime},m^{\prime}\rangle\propto\langle l,m|r=1,q=0;l^{\prime},m^{\prime}\rangle\langle\alpha;l||\hat{\mathcal{\mathbf{O}}}^{(1)}||\beta;l^{\prime}\rangle,$
with the first term being a Clebsch-Gordan coefficient and the second the reduced matrix element for tensor $\hat{\mathcal{\mathbf{O}}}^{(1)}=\Bigl{(}\hat{\mathcal{O}}_{+1}^{(1)},\hat{\mathcal{O}}_{0}^{(1)},\hat{\mathcal{O}}_{-1}^{(1)}\Bigr{)}$. Put it another way, the theorem states that operating with a rank-$r$ spherical tensor with states possessing an angular momentum $l$ is equivalent to studying the composition of two angular momenta, one with spin $r$ and the other with spin $l$. This analogy allows to define selection rules for the Clebsch-Gordan coefficients, which for the case at study translate in:
$\begin{cases}|l-r|\leq l^{\prime}\leq l+r\Rightarrow\Delta l=0,\pm 1\\ m^{\prime}=m+q\Rightarrow m^{\prime}=m\\ \end{cases}\$ (7)
@@
This property leads to the following key result: given an initial state with $b$ excitations, i.e. $b$ spin-up spins oriented along $z$ and so total magnetization $m=-L/2+b$; this initial state has support exclusively on the $b+1$ eigenspaces $l=L/2,L/2-1,\dots,L/2-b$, thus, due the selection rules in Eq.(7), we can observe exactly $b$ frequencies in the power spectrum of $\langle\hat{\mathcal{S}}^{z}_{i}\rangle$.
@@
Furthermore, the selection rule $m^{\prime}=m$, valid for $q=0$, allows to distinguish between the different subspaces of the total $z$-magnetization ^$\hat{\mathcal{M}}_{z}$ in case where the initial state is a superposition of states belonging to diverse magnetization sectors, like in Fig. 1$(b)$. Considering this example, the initial state lives among the levels $(l=L/2,m=-L/2),(l=L/2,m=-L/2+1),(l=L/2-1,m=-L/2+1)$; the described selection rule isolates the total $z$-magnetization sector between them, so that the frequencies originated from the mechanism above for $m=-L/2+1$ do not mix up with those arising from the subspace with $m=-L/2$, with this remaining true for more general situations.
@@
As a further numerical demonstration, we dub $d_{mn}$ the coefficients depending on the eigenbasis overlaps with the initial state and the operator’s matrix elements. The weights on the different exponentials are given by the sums of the corresponding $|d_{mn}|^{2}$ coefficients. We show in Fig. 3(b) the numerical evaluation of the coefficients normalized to their total sum. We notice that the position of the two resulting finite frequencies perfectly matches that displayed in the numerically simulated power spectrum, Fig.3(a). In particular, we observe a zero weight for $\Delta E_{31}$, in agreement with the absence of this frequency in the simulated spectrum. While the value of the frequencies depends on the system size $L$, we show in Fig. 3(c) that their number does not. In fact, the number of frequencies only depends on the number of initial excitations, as it is visible in Fig. 3(d) for the case of 3 initial excitations, where one additional finite frequency adds on.
@@
This analytical result embodies the physical interpretation of the simulational findings from [YagoMalo2023, yagomalo2024review]: the all-to-all connected quantum network is capable of preserving long-lasting memory in space and time of the number of initially injected excitations, this memory being encoded in frequency spectrum of the collective spin excitations.
@@
Figure 3: Multiple excitations with all-to-all connected XX network: frequency analysis. $(a)$ Spectrum’s peaks for the local $z$-magnetization signal in the case of two excitations: the corresponding positions are $5.02\pm 0.07$ (higher prominence) and $7.02\pm 0.08$ (lower prominence). The system parameters are the same as in Fig.1, with the signal being picked up from site 4. $(b)-(d)$ Bar plot representing the distribution of the coefficients $|d_{mn}|^{2}$ in eq. (6), showing that the number of frequencies perfectly matches with the simulated spectra in Fig.1. As in the case of the spectra, the zero-frequency is not shown to highlight the oscillating components of the signals. The coefficients are summed over the levels’ degeneracies and normalized to the total sum of the coefficients. $(b)$ $L=7$ and two excitations. $(c)$ $L=7$ and three excitations. $(d)$ $L=9$ and four excitations.

### v2

@@ 2.1 Eigenspaces of the exchange (XX) Hamiltonian
We start by analyzing only the exchange part of the Hamiltonian (1) with all-to-all interactions, and considering it as our unperturbed mean-field component:
$\hat{\mathcal{H}}_{0}=J\sum_{i<j}(\hat{\mathcal{S}}_{i}^{x}\hat{\mathcal{S}}_{j}^{x}+\hat{\mathcal{S}}_{i}^{y}\hat{\mathcal{S}}_{j}^{y})=\frac{J}{2}\sum_{i<j}(\hat{\mathcal{S}}_{i}^{+}\hat{\mathcal{S}}_{j}^{-}+\hat{\mathcal{S}}_{i}^{-}\hat{\mathcal{S}}_{j}^{+})\,.$
We show below that $\hat{\mathcal{H}}_{0}$ can be rewritten as function of the total $z$-magnetization $\hat{\mathcal{M}}_{z}=\sum_{i}\hat{\mathcal{S}}_{i}^{z}$ and of the (Casimir invariant) total squared magnetization $\hat{\mathcal{M}}^{2}=\hat{\mathcal{M}}_{x}^{2}+\hat{\mathcal{M}}_{y}^{2}+\hat{\mathcal{M}}_{z}^{2}$, see [70]. We then re-cast the operators $\hat{\mathcal{S}}_{i}^{x}\hat{\mathcal{S}}_{j}^{x}$ (equivalent for the $y$-component) in terms of $\hat{\mathcal{M}}_{z}$ and $\hat{\mathcal{M}}^{2}$ as:
$\hat{\mathcal{M}}_{x}^{2}=\sum_{k}(\hat{\mathcal{S}}_{k}^{x})^{2}+2\sum_{i<j}\hat{\mathcal{S}}_{i}^{x}\hat{\mathcal{S}}_{j}^{x}\,\Leftrightarrow\,\sum_{i<j}\hat{\mathcal{S}}_{i}^{x}\hat{\mathcal{S}}_{j}^{x}=\frac{\hat{\mathcal{M}}_{x}^{2}-\sum_{k}(\hat{\mathcal{S}}_{k}^{x})^{2}}{2}\,.$
@@
$\hat{\mathcal{H}}_{0}=\frac{J}{4}\Bigl{[}2(\hat{\mathcal{M}}^{2}-\hat{\mathcal{M}}_{z}^{2})-L\Bigr{]},$ (2)
with $L$ the system size (number of spins). The result in Eq. (2) implies that although the starting Hamiltonian is the sum of two-body interacting terms, the all-to-all connectivity allows for the appearance of collective behavior, where the system behaves as if it had a single, large spin [75]. Labeling the quantum numbers associated to operators $\hat{\mathcal{M}}^{2}$ and $\hat{\mathcal{M}}_{z}$ with $l$ and $m$, respectively, we use the angular momenta composition rules for spin-$1/2$ to calculate the corresponding energy bands $E_{lm}$:
[…]
$m$ $=-l,(-l+1),\dots,(l-1),l\quad\text{($2l+1$ values)}$
@@
Degeneracy-wise, we can compute the number $f(l,L)$ of different ways one can get a certain value of l for a given $L$, through the following recursive formula:
@@
$\text{deg}(E_{l|m|},L)=\begin{cases}2f(l,L)\quad|m|\neq 0\\ f(l,L)\quad|m|=0\\ \end{cases}\,.$
We give a proof and details of the computation in Appendix A.
Figure 2: Symmetries and structure of the energy levels for the all-to-all connected $X$ Hamiltonian. $(a)$ Diagram for a system with $L=7$ sites and constant exchange coupling $J/2=1$ (unperturbed study). The vertical axis represents the exact values of the eigenvalues of $\hat{\mathcal{H}}_{0}$, while the numbers on the left of the horizontal lines account for the degeneracy of the corresponding level; different colors and symbols represent different subspaces corresponding to different number of excitations along the $z$-axis. Red triangles and lines: levels corresponding to 1 excitation. Blue lozenges and lines: levels referring to 2 excitations. Double-headed arrows : the three possible energy gaps described by Eq. (5): the dotted arrow represents the frequency that is not manifested in the $z$-magnetization spectrum. Green bars: remaining energy levels. $(b)$ The same diagram for the case $L=9$, with the red bars contoured by stars representing the levels with 4 excitations denoted by the quantum number $l$ on the right.
@@
We are now in a position to explain the local $z$-magnetization spectra in Fig. 1 in terms of the energy levels. Due to the $U(1)$ symmetry, the quantum number $m$ will be fixed during the time evolution. If we consider the case of one single excitation to start with, we have $m_{1}=(-L/2)+1$, hence only the values $l_{1}=(L/2)-1$ and $l_{2}=L/2$ are allowed, the corresponding bands having energies $E_{l_{1}m_{1}}=-{J}/2\doteq E_{1}^{(1)}$ with degeneracy L1L-1 and $E_{l_{2}m_{1}}={J(L-1)}/2\doteq E_{2}^{(1)}$ with degeneracy 1. Thus, the oscillation of the local $z$-magnetization signal occurs at angular frequency equal to the energy split between the two levels:
@@
We notice that the local $z$-magnetization spectra depend explicitly on the separation between the energy levels of the system, see vertical dashed line in Fig. 1(c)(c).
We can use the same line of reasoning for the second case highlighted in Fig. 2(a)(a), with two excitations. Here, the quantum number $m$ is fixed to be $m_{2}=(-L/2)+2$, implying that $l$ can take values $l_{1}=(L/2)-2$, $l_{2}=(L/2)-1$ and $l_{3}=L/2$, reflecting in the energy levels as $E_{l_{1}m_{2}}=-J\doteq E_{1}^{(2)}$ with degeneracy L(L3)/2L(L-3)/2; $E_{l_{2}m_{2}}={J(L-4)}/2\doteq E_{2}^{(2)}\,\text{with degeneracy $L-1$;}$ $E_{l_{3}m_{2}}=J(L-2)\doteq E_{3}^{(3)}\,\text{ with degeneracy 1.}$ The resulting energy gaps are:
@@
Not all the possible energy differences appear in the power spectrum of the dynamics. To show this, we analyze in Fig. 3 the power spectrum of the local magnetization $\langle\hat{\mathcal{S}}_{i}^{z}\rangle(t)$. For instance, in the case of two excitations, see Fig. 3(a), we observe the presence of two frequencies, corresponding to the gaps $\Delta E_{21}$ and $\Delta E_{32}$, while the third possible transition is absent. Even in the case depicted in Fig.1(d)(e)(d)-(e), corresponding to four initial excitations or m=L/2+4m=-L/2+4, we ascertain the presence of four out of the $4\cdot(4-1)/2=6$ possible frequencies arising from the five levels having quantum number l=L/2,L/21,L/22,L/23,L/24l=L/2,L/2-1,L/2-2,L/2-3,L/2-4; again, these four frequencies are exactly those coming from the difference between contiguous levels. The schematics for this situation is shown in Fig. 2(b)(b).
@@
where we have decomposed the initial state $|\psi_{0}\rangle$ temporal evolution in the Hamiltonian eigenbasis $\{|\phi_{n}\rangle\}$, with corresponding eigenvalues $\{E_{n}\}$ (the pedices here represent generic quantum numbers). In our specific case, the eigenstates of the system Hamiltonian $\hat{\mathcal{H}}_{0}$ are those of the angular momentum $|\alpha;l,m\rangle$, with $l$ and $m$ defined in Eq. 3 and $\alpha$ another quantum number independent of angular momentum, the initial state is a superposition of Pauli strings, and the operator $\hat{\mathcal{O}}$ is a local spin operator $\hat{\mathcal{S}}^{z}_{i}\doteq\hat{\mathcal{O}}_{0}^{(1)}$ at site $i$: this operator has the property of being the $q=0$ component of a spherical tensor of rank $r=1$, the other components being $\hat{\mathcal{O}}_{+1}^{(1)}=-c\hat{\mathcal{S}}^{+}_{i}$ and $\hat{\mathcal{O}}_{-1}^{(1)}=c\hat{\mathcal{S}}^{-}_{i}$ with $c$ a positive constant. The coefficients $\langle\psi_{0}|\phi_{m}\rangle=\langle\psi_{0}|\alpha;l,m\rangle$ and $\langle\phi_{n}|\psi_{0}\rangle=\langle\beta;l^{\prime},m^{\prime}|\psi_{0}\rangle$ are the Clebsch-Gordan coefficients between the eigenbasis and the computational basis. The matrix elements of $\hat{\mathcal{O}}_{0}^{(1)}$ can be expressed thanks to the Wigner-Eckart theorem [76] as:
[…]
with the first term being a Clebsch-Gordan coefficient and the second the reduced matrix element for tensor $\hat{\mathcal{\mathbf{O}}}^{(1)}=\Bigl{(}\hat{\mathcal{O}}_{+1}^{(1)},\hat{\mathcal{O}}_{0}^{(1)},\hat{\mathcal{O}}_{-1}^{(1)}\Bigr{)}$. Put it another way, the theorem states that operating with a rank-$r$ spherical tensor with states possessing an angular momentum $l$ is equivalent to studying the composition of two angular momenta, one with spin $r$ and the other with spin $l$. This analogy allows to define selection rules for the Clebsch-Gordan coefficients, which for the case at study translates in:
[…]
@@
This property leads to the following key result: given an initial state with $b$ excitations, i.e. $b$ spin-up spins oriented along $z$ and so total magnetization m=L/2+bm=-L/2+b; this initial state has support exclusively on the $b+1$ eigenspaces $l=L/2,L/2-1,\dots,L/2-b$, thus, due the selection rules in Eq.(7), we can observe exactly $b$ frequencies in the power spectrum of $\langle\hat{\mathcal{S}}^{z}_{i}\rangle$.
@@
Furthermore, the selection rule $m^{\prime}=m$, valid for $q=0$, allows to distinguish between the different subspaces of the total $z$-magnetization $\hat{\mathcal{M}}_{z}$ in case where the initial state is a superposition of states belonging to diverse magnetization sectors, like in Fig. 1(b)(b). Considering this example, the initial state lives among the levels (l=L/2,m=L/2),(l=L/2,m=L/2+1),(l=L/21,m=L/2+1)(l=L/2,m=-L/2),(l=L/2,m=-L/2+1),(l=L/2-1,m=-L/2+1); the described selection rule isolates the total $z$-magnetization sector between them, so that the frequencies originated from the mechanism above for m=L/2+1m=-L/2+1 do not mix up with those arising from the subspace with m=L/2m=-L/2, with this remaining true for more general situations.
@@
As a further numerical demonstration, we dub $d_{mn}$ the coefficients depending on the eigenbasis overlaps with the initial state and the operator’s matrix elements, see Eq. (6). The weights on the different exponentials are given by the sums of the corresponding $|d_{mn}|^{2}$ coefficients. We show in Fig. 3(b) the numerical evaluation of the coefficients normalized to their total sum. We notice that the position of the two resulting finite frequencies perfectly matches that displayed in the numerically simulated power spectrum, Fig.3(a). In particular, we observe a zero weight for $\Delta E_{31}$, in agreement with the absence of this frequency in the simulated spectrum. While the value of the frequencies depends on the system size $L$, we show in Fig. 3(c) that their number does not. In fact, the number of frequencies only depends on the number of initial excitations, as it is visible in Fig. 3(d) for the case of 3 initial excitations, where one additional finite frequency adds on.
@@
This analytical result embodies the physical interpretation of the simulational findings from [5, 14]: the all-to-all connected quantum network is capable of preserving long-lasting memory in space and time of the number of initially injected excitations, this memory being encoded in frequency spectrum of the collective spin excitations.
@@
Figure 3: Multiple excitations with all-to-all connected $X$ network: frequency analysis. $(a)$ Spectrum’s peaks for the local $z$-magnetization signal in the case of two excitations: the corresponding positions are 5 (higher prominence) and 7 (lower prominence). The system parameters are the same as in Fig.1, with the signal being picked up from site 4. (b)(d)(b)-(d) Bar plot representing the distribution of the coefficients $|d_{mn}|^{2}$ in eq. (6), showing that the number of frequencies perfectly matches with the simulated spectra in Fig.1. As in the case of the spectra, the zero-frequency is not shown to highlight the oscillating components of the signals. The coefficients are summed over the levels’ degeneracies and normalized to the total sum of the coefficients. $(b)$ $L=7$ and two excitations. $(c)$ $L=7$ and three excitations. $(d)$ $L=9$ and four excitations.